# Agentic Conversational Search with Contextualized Reasoning via Reinforcement Learning

### v1

@@ 1 Introduction
Large Language Models (LLMs) have become a popular interface for human–AI interaction, supporting information seeking and task completion assistance through natural, multi-turn dialogue Gao et al. (2022); Zamani et al. (2023); Mo et al. (2024b). To respond to users within multi-turn dialogues, the context-dependent user intent usually evolves across historical interactions, requiring contextual interpretation Jin et al. (2023), mixed-initiative actions Aliannejadi et al. (2021), and dynamic coordination between information retrieval and response generation Qian and Liu (2025).
@@
To facilitate the multi-turn conversational scenarios, existing studies Roy et al. (2024); Ye et al. (2024); Mo et al. (2026a) usually follow static “rewrite, retrieve, and generate” workflows with separated models. Such a pipeline optimizes various procedures separately and requires different task-specific supervision signals, including rewritten queries, relevance judgments, and ground-truth answers for each turn. Besides, it might overlook the mixed-initiative action optimization, e.g., asking a clarification question at a suitable moment, simultaneously. Aiming for a dynamic procedure, recent developments in deep search agentic models Jin et al. (2025); Zheng et al. (2025); Song et al. (2025) achieve remarkable performance on complex information-seeking tasks by interleaving multi-round reasoning and external search. The flexible architecture allows it to dynamically explore external knowledge sources and iteratively refine its reasoning trajectory for open-domain question answering Zhang et al. (2024b); Zhu et al. (2025b); Zhang et al. (2025a). However, these advanced deep search agentic models are designed for single-turn information accessing tasks, which might lack the ability to handle multi-turn conversational scenarios Zhang et al. (2024a); Zhu et al. (2025a).
@@
Under user-system multiple interactions, the user’s queries are context-dependent and involve topic-switch Mao et al. (2022b); Mo et al. (2023b); Yoon et al. (2025). Thus, it is crucial for the conversational agent to learn to leverage useful historical context to reply to the current turn via reasoning ability. One important aspect is to generate a stand-alone search query condition on comprehending the historical context Jang et al. (2023); Mo et al. (2024a), which is different from the existing single-turn deep search agent that decomposes a self-contained query. Another crucial factor lies in optimizing the model to effectively leverage information obtained from search results through de-contextualization, enabling accurate answer generation that integrates retrieved knowledge with the model’s inherent parametric knowledge Wang et al. (2024a); Su et al. (2025); Mo et al. (2026b), while the existing studies on deep search lack search results optimization Jin et al. (2025). Besides, the flexible interaction interface enables the system to respond in various ways, which make it possible to equip the conversational agents with mixed-initiative actions capabilities to response in suitable forms, such as asking clarification questions Zou et al. (2023), providing short or long form answers with dynamic decisions depending on the query type Su et al. (2024); Mo et al. (2025b), and rejecting to answer when without enough evidence Wang et al. (2024b). The mixed-initiative actions are useful for the system to engage with the user Aliannejadi et al. (2021); Wang et al. (2024c).
@@
To address the aforementioned issues, we aim to optimize the agentic conversational search task from multiple aspects within a single agentic framework with contextualized reasoning, including history-conditional search query generation, the utilization of search results for accurate answer generation, and interactive action decision via mixed-initiative capacities. To achieve it, we decompose the overall reward into three complementary components under a reinforcement learning (RL) training framework Shao et al. (2024) to interleave search and reasoning across turns, enabling exploratory and adaptive behaviors learned through tailored design. Specifically, to help the agentic model optimize search queries through de-contextualization reasoning in conversations and learn to extract useful information from retrieved results for answer generation, we introduce a search optimization reward. This reward measures the information overlap between the retrieved content and the ground-truth answers, providing additional supervision beyond the ground-truth answer signal alone. Besides, a mixed-initiative action reward is used to enhance the alignment with evolving user goals and facilitate the mixed-initiative capacity of the models. The experimental results across four widely used conversational search benchmarks demonstrate the effectiveness of our methods by surpassing several existing strong baselines with different principles under various settings. Detailed ablation analyses show the contributions of different components and insightful observations to improve multi-turn capacity for conversational agents.
@@ 2 Related Work
Conversational Question Answering is the task of responding to user queries within multi-turn dialogues, with dependence on previous turns and often grounding from external knowledge Qu et al. (2020). The retrieval-augmented generation (RAG) technique Lewis et al. (2020) is a common practice to obtain external factual grounding in answers. Previous studies in conversational search Wu et al. (2022); Mo et al. (2023a); Abbasiantaeb et al. (2024); Lupart et al. (2025b) aim to enhance the understanding of user intent and more complex user interactions. A series of approaches are proposed to improve the search results performance in conversations, including query rewriting Ye et al. (2023); Lai et al. (2024), conversational dense retrieval fine-tuning Lin et al. (2021); Kim and Kim (2022); Mao et al. (2024), and data augmentation Mao et al. (2022a); Mo et al. (2025c). Besides, several studies Meng et al. (2023); Owoicho et al. (2023) have explored improving mixed-initiative action detection in conversational settings to provide more effective support during interactions. However, these approaches do not directly explore how to leverage the retrieved information to improve answer generation in the era of LLMs Gao et al. (2023); Mo et al. (2024b); Bhaskar et al. (2025); Zhang et al. (2025b, c). Our research aims to develop an agentic conversational search model that can understand context-dependent user intent, optimize search queries, enhance mixed-initiative interaction, and reason over search results to generate accurate answers.
@@ 3.2 Rewards Modeling
Reinforcement learning (RL) plays a crucial role in training agentic models by enabling self-exploration of diverse reasoning trajectories and rewarding those that are effective. This aligns model behavior with the objectives of information seeking for users’ context-dependent intents and efficient decision-making for actions in conversations. However, since these trajectories are self-generated and do not always yield correct final answers due to noisy conversational context, relying solely on outcome-based rewards yields sparse supervision signals, making the training process harder to optimize Chen et al. (2025). To this end, we decompose the overall reward into three complementary components: outcome reward, search optimization reward, and mixed-initiative action reward. These rewards guide the model to reason more strategically with the conversational context towards the final answer, effectively capture user intent for search optimization, and respond precisely to various types of queries with mixed-initiative needs in conversations. We describe the overview of our framework in Figure 1.
@@ 3.2.3 Mixed-Initiative Action Reward
The mixed-initiative plays an important role in conversational scenarios, where the agentic model is expected to clarify ambiguous query turns or refuse to answer when their knowledge is insufficient. To facilitate the mixed-initiative capacity of our model, we formulate the mixed-initiative as a classification task to validate whether the agentic model can decide the necessary supportive reactions. Specifically, we include four types of mixed-initiative action following Wu et al. (2023) in the reaction set as $\mathcal{A}\text{ct}=\{\text{answer},\text{clarify},\text{no answer}$. In addition to generating a final answer as a common case, the agentic model should generate clarification questions, return only relevant search results with the final answer, or refuse to answer to accommodate the unanswerable cases. Then, the mixed-initiative action (MIA) reward $\mathcal{R}_{\text{MIA}}$ is designed similarly to the format reward to measure whether the specific label tokens, i.e., <clarify> and <noanswer>, are included in the generated sequence $r_{n}$ among the turns that are annotated with requiring specific mixed-initiative actions in the ground-truth reaction set $\mathcal{A}\text{ct}_{n}^{*}$ as
@@ 4.3 Ablation Studies
Ours w/o $\mathcal{R}_{\text{action}}$ (Base-7b) 1.42 3.67 - -
Ours w/. $\mathcal{R}_{\text{action}}$ & sep. label 11.51 83.19 3.9 5.8
Ours w/. $\mathcal{R}_{\text{action}}$ & comb. label 6.99 78.35 0.2 0.8

### v2

@@ 1 Introduction
Large Language Models (LLMs) have become a popular interface for human–AI interaction, supporting information seeking and task completion assistance through natural, multi-turn dialogue Gao et al. (2022); Zamani et al. (2023); Mo et al. (2024b). To respond to users within multi-turn dialogues, the context-dependent user intent usually evolves across historical interactions, requiring contextual interpretation Jin et al. (2023), mixed-initiative actions Aliannejadi et al. (2021), and dynamic coordination between information retrieval and response generation Su et al. (2024); Qian and Liu (2025); Lai et al. (2025); Qi et al. (2026).
@@
To facilitate the multi-turn conversational scenarios, existing studies Roy et al. (2024); Ye et al. (2024); Mo et al. (2026a) usually follow static “rewrite, retrieve, and generate” workflows with separated models. Such a pipeline optimizes various procedures separately and requires different task-specific supervision signals, including rewritten queries, relevance judgments, and ground-truth answers for each turn. Besides, it might overlook the mixed-initiative action optimization, e.g., asking a clarification question at a suitable moment, simultaneously. Aiming for a dynamic procedure, recent developments in deep search agentic models Jin et al. (2025); Zheng et al. (2025); Zeng et al. (2026) achieve remarkable performance on complex information-seeking tasks by interleaving multi-round reasoning and external search. The flexible architecture allows it to dynamically explore external knowledge sources and iteratively refine its reasoning trajectory for open-domain question answering Zhang et al. (2024b); Zhu et al. (2025b); Zhang et al. (2025a); Su et al. (2025). However, these advanced deep search agentic models are designed for single-turn information accessing tasks, which might lack the ability to handle multi-turn conversational scenarios Zhang et al. (2024a); Zhu et al. (2025a); Mo et al. (2025c).
@@
Under user-system multiple interactions, the user’s queries are context-dependent and involve topic-switch Mao et al. (2022b); Mo et al. (2023b); Yoon et al. (2025). Thus, it is crucial for the conversational agent to learn to leverage useful historical context to reply to the current turn via reasoning ability. One important aspect is to generate a stand-alone search query condition on comprehending the historical context Jang et al. (2023); Mo et al. (2024a), which is different from the existing single-turn deep search agent that decomposes a self-contained query. Another crucial factor lies in optimizing the model to effectively leverage information obtained from search results through de-contextualization, enabling accurate answer generation that integrates retrieved knowledge with the model’s inherent parametric knowledge Wang et al. (2024a); Su et al. (2025); Mo et al. (2026b), while the existing studies on deep search lack search results optimization Jin et al. (2025). Besides, the flexible interaction interface enables the system to respond in various ways, which make it possible to equip the conversational agents with mixed-initiative actions capabilities to response in suitable forms, such as asking clarification questions Zou et al. (2023), providing short or long form answers with dynamic decisions depending on the query type Jeong et al. (2024); Mo et al. (2025b), and rejecting to answer when without enough evidence Wang et al. (2024b). The mixed-initiative actions are useful for the system to engage with the user Aliannejadi et al. (2021); Mo et al. (2024d); Wang et al. (2024c).
@@
To address the aforementioned issues, we aim to optimize the agentic conversational search task from multiple aspects within a single agentic framework with contextualized reasoning, including history-conditional search query generation, the utilization of search results for accurate answer generation, and interactive action decision via mixed-initiative capacities. To achieve it, we propose ConvAgent to decompose the overall reward into three complementary components under a reinforcement learning (RL) training framework Shao et al. (2024) to interleave search and reasoning across turns, enabling exploratory and adaptive behaviors learned through tailored design. Specifically, to help the agentic model optimize search queries through de-contextualization reasoning in conversations and learn to extract useful information from retrieved results for answer generation, we introduce a search optimization reward. This reward measures the information overlap between the retrieved content and the ground-truth answers, providing additional supervision beyond the ground-truth answer signal alone. Besides, a mixed-initiative action reward is used to enhance the alignment with evolving user goals and facilitate the mixed-initiative capacity of the models. The experimental results across four widely used conversational search benchmarks demonstrate the effectiveness of our methods by surpassing several existing strong baselines with different principles under various settings. Detailed ablation analyses show the contributions of different components and insightful observations to improve multi-turn capacity for conversational agents.
@@ 2 Related Work
Conversational Question Answering is the task of responding to user queries within multi-turn dialogues, with dependence on previous turns and often grounding from external knowledge Qu et al. (2020). The retrieval-augmented generation (RAG) technique Lewis et al. (2020) is a common practice to obtain external factual grounding in answers. Previous studies in conversational search Wu et al. (2022); Mo et al. (2023a); Abbasiantaeb et al. (2024); Lupart et al. (2025b) aim to enhance the understanding of user intent and more complex user interactions. A series of approaches are proposed to improve the search results performance in conversations, including query rewriting Mao et al. (2023); Ye et al. (2023); Lai et al. (2024), conversational dense retrieval fine-tuning Lin et al. (2021); Kim and Kim (2022); Mao et al. (2024); Mo et al. (2024c), and data augmentation Mao et al. (2022a); Mo et al. (2025d). Besides, several studies Meng et al. (2023); Owoicho et al. (2023) have explored improving mixed-initiative action detection in conversational settings to provide more effective support during interactions. However, these approaches do not directly explore how to leverage the retrieved information to improve answer generation in the era of LLMs Gao et al. (2023); Mo et al. (2024b); Bhaskar et al. (2025); Zhang et al. (2025b, c). Our research aims to develop an agentic conversational search model that can understand context-dependent user intent, optimize search queries, enhance mixed-initiative interaction, and reason over search results to generate accurate answers.
@@ 3.2 Rewards Modeling
Reinforcement learning (RL) plays a crucial role in training agentic models by enabling self-exploration of diverse reasoning trajectories and rewarding those that are effective. This aligns model behavior with the objectives of information seeking for users’ context-dependent intents and efficient decision-making for actions in conversations. However, since these trajectories are self-generated and do not always yield correct final answers due to noisy conversational context, relying solely on outcome-based rewards yields sparse supervision signals, making the training process harder to optimize Chen et al. (2025); Zhang et al. (2026). To this end, we decompose the overall reward into three complementary components: outcome reward, search optimization reward, and mixed-initiative action reward. These rewards guide the model to reason more strategically with the conversational context towards the final answer, effectively capture user intent for search optimization, and respond precisely to various types of queries with mixed-initiative needs in conversations. We describe the overview of our framework in Figure 1.
@@ 3.2.3 Mixed-Initiative Action Reward
The mixed-initiative plays an important role in conversational scenarios, where the agentic model is expected to clarify ambiguous query turns or refuse to answer when their knowledge is insufficient. To facilitate the mixed-initiative capacity of our model, we formulate the mixed-initiative as a classification task to validate whether the agentic model can decide the necessary supportive reactions. Specifically, we include four types of mixed-initiative action following Wu et al. (2023) in the reaction set as $\mathcal{A}\text{ct}=\{\text{answer},\text{clarify},\text{no answer}\}$. In addition to generating a final answer as a common case, the agentic model should generate clarification questions, return only relevant search results with the final answer, or refuse to answer to accommodate the unanswerable cases. Then, the mixed-initiative action (MIA) reward $\mathcal{R}_{\text{MIA}}$ is designed similarly to the format reward to measure whether the specific label tokens, i.e., <clarify> and <noanswer>, are included in the generated sequence $r_{n}$ among the turns that are annotated with requiring specific mixed-initiative actions in the ground-truth reaction set $\mathcal{A}\text{ct}_{n}^{*}$ as
@@ 4.2.3 Correlation between Search Quality and Answer Generation Performance
ConvAgent w/o $\mathcal{R}_{\text{action}}$ (Base-7b) 1.42 3.67 - -
ConvAgent w/. $\mathcal{R}_{\text{action}}$ & sep. label 11.51 83.19 3.9 5.8
ConvAgent w/. $\mathcal{R}_{\text{action}}$ & comb. label 6.99 78.35 0.2 0.8